# ReassembleNet: Learnable Keypoints and Diffusion for 2D Fresco Reconstruction

### v2

@@ Abstract
The task of reassembly is a significant challenge across multiple domains, including archaeology, genomics, and molecular docking, requiring the precise placement and orientation of elements to reconstruct an original structure. In this work, we address key limitations in state-of-the-art Deep Learning methods for reassembly, namely i) scalability; ii) multimodality; and iii) real-world applicability: beyond square or simple geometric shapes, realistic and complex erosion, or other real-world problems. We propose ReassembleNet, a method that reduces complexity by representing each input piece as a set of contour keypoints and learning to select the most informative ones by Graph Neural Networks pooling inspired techniques. ReassembleNet effectively lowers computational complexity while enabling the integration of features from multiple modalities, including both geometric and texture data. Further enhanced through pretraining on a semi-synthetic dataset. We then apply diffusion-based pose estimation to recover the original structure. We improve on prior methods by $55\%$ and $86\%$ for RMSE Rotation and Translation, respectively.
@@ 1 Introduction
Reassembly requires placing each element in its correct position and orientation to form the original shape as a whole – whether it be a 2D or 3D object. This ability is a form of spatial intelligence, which refers to the capacity to accurately perceive the visual-spatial environment and manipulate that perceived space [9]. This skill is typically evaluated through reassembly tasks, where individual components must be arranged and connected to form a coherent, functional whole—such as solving 2D jigsaw puzzles or assembling 3D structures with LEGO blocks.
Since the advent of the first puzzle solver [5], reassembly tasks have posed a significant challenge to the machine learning community due to their inherent combinatorial complexity. These challenges are further underscored by their wide range of applications, including genomics [21], assistive technologies [39], and molecular docking [3].
@@
In recent years, AI techniques for reassembling objects have gained increasing traction in the heritage field, particularly in fresco reconstruction [38]. This growing application of AI is driven by the fundamental challenge archaeologists face in reconstructing the past. After extensive and painstaking work involving site surveys and excavations, they are often confronted with the daunting task of reassembling countless fragments of varying sizes, shapes, and appearances to recreate ancient artifacts or artworks. Depending on the complexity of the artifact, this reassembly process can span months, years, or even decades. In cases involving a vast number of pieces, the task may become nearly insurmountable, regardless of the skill of the experts. Despite recent advances, state-of-the-art methods for solving this task remain far from achieving a satisfactory accuracy [36]. The challenges persist not only in the 3D domain but also in the 2D space, where the reduced degrees of freedom do not significantly alleviate the complexity of the problem. Reduction to 2D has been used for a long time to aid in finding the solutions as many heritage objects, e.g. frescos, are planar on the dominant surface [6]. Traditional approaches use keypoints or patches [25, 22, 35, 4] benefiting from optimization strategy allowing them to scale to complex irregular shapes, in contrast to deep learning approaches. Three primary factors contribute to the difficulty of fresco reassembly for deep learning approaches. First, the inherent complexity of irregular pieces and the diverse textures present in the fragments (Figure 1). Secondly, as a byproduct of the first, complex polygonal shapes are difficult to model in deep learning approaches, where comparisons need to be made across the points of one fragment edge to another fragment. Finally, deep learning methods typically require large-scale datasets to effectively learn meaningful patterns [32], a requirement that is difficult to meet in most applications domains of reassembly.
@@
To address these challenges, we propose ReassembleNet, a scalable deep learning method that integrates a diffusion process with multi-modal feature fusion. Our approach enables scalability on irregular 2D shapes through keypoint selection, a crucial capability for highly fragmented structures that existing methods struggle to process [18]. We introduce a semi-synthetic dataset specifically designed to highlight the characteristics of fresco pieces. Through fine-tuning, we demonstrate that pre-training on a semi-synthetic dataset significantly enhances model performance on real-world frescos. We combine the keypoint selection with geometry and texture representation to constrain the matching problem modeled through an inter and intra-piece attention block. Finally, we wrap this model in a diffusion process to iteratively denoise the pose of pieces into the final locations.
@@ 2.1 Reassembly of Square-based puzzles
Square-based jigsaws have been extensively studied in the literature, with early approaches using greedy methods. For instance, [24] introduced a three-phase greedy approach involving placement, segmentation, and shifting, relying on compatibility functions and estimation metrics. While effective for large puzzles, it requires predefined orientations and does not handle missing pieces. Similarly, [7] proposed a tree-based algorithm for puzzles with unknown orientations and locations, using Mahalanobis Gradient Compatibility (MGC) for boundary analysis. However, it struggles with color-based similarity and missing pieces. [23] presented a greedy strategy for puzzles without prior information on piece sizes or orientations, refining the initial configuration. It works with mixed or missing pieces but struggles with noisy images. Other methods, such as [19], combine edge and content similarity using MGC, but face challenges with large-scale puzzles. [29] used a genetic algorithm based on color similarity to solve puzzles with unknown locations and orientations.
In recent years, several deep learning approaches have been developed for puzzle reconstruction. One prominent direction leverages generative models to reconstruct a complete image from an unordered set of puzzle pieces [33, 34]. These methods aim to infer the global structure of the puzzle, effectively generating missing spatial relationships. However, they are not applicable when the input pieces are rotated, as they assume a fixed orientation. Another line of research explores diffusion-based processes to iteratively refine the placement of puzzle pieces step by step [10, 26]. These approaches have demonstrated effectiveness in solving the oriented puzzle problem, where piece orientations are known. However, they are restricted to regular puzzle settings, where pieces conform to a structured grid. For real-world applications, square-based approaches are often impractical, as broken objects typically have irregular shapes, noisy visual content, and eroded or missing edges between pieces. A naive solution is to create a regular patch that encloses the broken object; however, this drastically hinders the effectiveness.
@@ 2.2 Reassembly of irregular pieces
Before the adoption of deep learning, various techniques were developed to address this problem. For instance, [25] proposed an artifact reassembly method using inpainting, texture synthesis, and FFT-based image registration. Their approach aligns fragments by maximizing correlation through FFT shift theory, effectively handling damaged edges. Similarly, [22] uses a divide-and-conquer strategy, grouping pieces based on texture and color similarity. The method utilizes RGB and HIS color spaces for color matching and employs co-occurrence matrices to extract texture features. In [35], the authors introduced a computer-aided method consisting of four key steps. Their approach reconstructs fragmented images by identifying adjacent pieces, matching contours using a Smith-Waterman-based method with color similarity, refining alignment through Iterative Closest Point, and assembling the final image based on optimized alignment angles. More recently, [4] proposed a patch-based optimization method for artifact reconstruction, leveraging color, gradients, and geometric transformations to match fragments.
In recent years, significant progress has been made in applying deep learning to handle irregular input shapes. The first work in this area is [30], which introduced DNN-Buddies, a deep neural network model for evaluating piece compatibility. In [20], the authors enhanced puzzle piece matching using CNNs and adaptive boosting. They proposed a multi-graph search algorithm to replace greedy strategies, enabling the handling of missing pieces and low-texture areas. [1] developed a classification network for evaluating the compatibility of fragment pairs and matching irregular puzzle pieces. However, their approach requires square-shaped pieces and cannot accommodate irregular outputs. PuzzleFusion [17] is based on a diffusion model and treats puzzle pieces as simple polygons with a limited number of vertices. This method is designed to work only on toy problems, as it processes all keypoints of the input polygons. As a result, it becomes impractical for real-world datasets, where a large number of keypoints lead to exponential memory consumption. Finally, PairingNet [40] is a learning-based image fragment pair-searching and -matching approach. Their method employs a graph-based network for feature extraction, a linear transformer-based module for fragment pair-searching, and a weighted fusion module for pair-matching. By formulating a similarity matrix to infer adjacent segments.
@@ 3 ReassembleNet Pipeline
In this paper, we focus on reassembling fragmented 2D objects where our method input is a set of images representing the fragments with irregular shapes. Following [18], our method takes as input $m$ images of unordered pieces. We extract keypoints from the images by applying Harris Corner Detector [13]. Points act as a helpful cue in the arrangement of the pieces as subsets of the points on each piece can be aligned to recover the overall piece pose.
@@
Next, we aim at extracting a multi-modal keypoint feature designed to capture both the geometric and textural properties of each fragment. This allows us to fully embed the shape and visual characteristics of the pieces, which are essential for proper alignment. Finally, we utilize a combination of attention blocks to model the relationships between the fragments. This solution helps the method to better fit the pieces together by analyzing their spatial and contextual connections, ensuring they are aligned in the correct configuration. The model is wrapped within a diffusion process to progressively learn the correct alignment in terms of translation ${\textbf{s}_{i}^{m}}\in~{}\mathbb{R}^{2}$ and rotation $\textbf{r}_{i}^{m}=[\cos(\theta_{i}^{m}),\sin(\theta_{i}^{m})]^{\top}$ [41].
@@ 3.1 KeyPoint Feature Representation
Texture Features. We have also focused on extracting features specifically related to the textures of the pieces. In particular, we extracted global texture features that summarize the general characteristics of the element, as well as local texture features that capture more detailed, localized information. For global texture features, we leveraged a pre-trained ResNet18 [15] to capture high-level texture representations across the entire element. For local texture features, we adopted a similar approach using the same pre-trained ResNet18. Specifically, we extract 32×32 patches centered on each selected keypoint. These localized patches are processed through ResNet18, where we remove the final classification layer and utilize the penultimate layer for feature extraction. This combination of global and local texture features allows our model to capture both the overall texture characteristics of the element as well as more detailed texture variations, which are crucial for accurately solving reassembly tasks.
@@ 3.2 Keypoint Selector Module
Non-Learnable Algorithm. For the non-learnable strategy, we employ the heuristic Farthest Point Sampling (FPS) [11] to perform an initial pre-selection of the $k$-points. FPS is a well-established technique in the field of 3D point cloud processing and geometric sampling. It iteratively selects the point that is furthest from the current set, ensuring maximum dispersion throughout the entire data space. This approach guarantees that the selected keypoints are evenly distributed, capturing the underlying structure and geometry of the input data effectively. By leveraging FPS, our method can robustly cover the entire image or spatial domain with a fixed number of representative points.
Learnable Algorithm. We frame the keypoint selection task as a graph sparsification problem [14]. In our formulation, the set of keypoints ${K}^{m}$ corresponds to the set of vertices $V^{m}$ of a complete graph $G^{m}$ (i.e. fully connected). Graph sparsification is then defined as the process of selecting a subset of nodes—i.e., keypoints—or edges from $G^{m}$ to produce a sparser graph $\hat{G}^{m}$. In other words, the elements of $\hat{G}^{m}$ are a subset of those in $G^{m}$.
@@
Graph Transformer. The projected features are then processed by a Graph Transformer [28] to aggregate and refine the input information.
Pooling Layer. Finally, we apply a pooling layer [8], where we select $k$ nodes from the original graph. The selection of nodes to drop is guided by a projection score computed against a learnable vector p. To ensure that gradients propagate into p, these projection scores also serve as gating values, allowing nodes with lower scores to retain fewer features. Defined as:
@@
[…]
@@ 3.3 Estimating position and rotation with diffusion
We adopt Diffusion Probabilistic Models as defined in Denoising Diffusion Implicit Models (DDIM) [31]. For each piece $m$, we apply an initial transformation, consisting of a translation $\textbf{s}^{m}$ and a rotation $\textbf{r}^{m}$, that is replicated identical for all its keypoints. To represent this initial transformation compactly, we define a concatenated vector for each keypoint $\textbf{x}_{i0}^{m}=[\textbf{s}_{i0}^{m^{\top}},\textbf{r}_{i0}^{m^{\top}}]^{\top}$, where $\textbf{s}^{m}_{i0}$ and $\textbf{r}^{m}_{i0}$ denote the initial translation and rotation applied to all keypoints of the piece. Since this transformation is applied globally, all keypoints within the same piece share these initial transformation parameters.
At training time, we iteratively add noise sampled from a Gaussian distribution $\mathcal{N}(0,I)$ to the poses of each keypoints (Forward Process). Following that, for the denoising process, we train ReassembleNet to reverse the noising process (Reverse Process) and predict the initial poses of all keypoints for every piece. We denote the initial poses of all the keypoints as $X_{0}=\{X_{0}^{m}\}_{m=1}^{M},$ where for each piece, $X^{m}_{0}=\{\textbf{x}^{m}_{i0}\}_{i=1}^{K},$ represents the set of initial keypoint poses. Additionally, we denote the predicted keypoint poses at timestep $t-1$ as $\hat{X}_{t-1}=\{\hat{X}_{t-1}^{m}\}_{m=1}^{M}$. For the final result at time $t=0$, we take the average polygon Euclidean center position and the polygon rotation. Further details on the diffusion process we use can be found in Supplementary Material C.
The Architecture. To process the task, we aim to capture both intra-piece information and inter-piece (i.e. global information between pieces). The key idea is to handle these two types of information separately but in parallel. To achieve this, we employ a combination of attention layers. Intra-piece information is processed using a sparse self-attention block [2], which helps the network focus on obtaining consistent positions and rotations, across different keypoints within the same piece. While, inter-piece information is handled using a standard self-attention block [37] between pieces.
Losses. Following [16] and standard practice in Diffusion Models, we train ReassembleNet to directly predict $\hat{X}_{0}$ rather than $\hat{X}_{t-1}$. We employ two loss functions to reconstruct the initial pose of each piece’s keypoints.

### v3

@@ Abstract
The task of reassembly is a significant challenge across multiple domains, including archaeology, genomics, and molecular docking, requiring the precise placement and orientation of elements to reconstruct an original structure. In this work, we address key limitations in state-of-the-art Deep Learning methods for reassembly, namely i) scalability; ii) multimodality; and iii) real-world applicability: beyond square or simple geometric shapes, realistic and complex erosion, or other real-world problems. We propose ReassembleNet, a method that reduces complexity by representing each input piece as a set of contour keypoints and learning to select the most informative ones by Graph Neural Networks pooling inspired techniques. ReassembleNet effectively lowers computational complexity while enabling the integration of features from multiple modalities, including both geometric and texture data. Further enhanced through pretraining on a semi-synthetic dataset. We then apply diffusion-based pose estimation to recover the original structure. We improve on prior methods by $57\%$ and $87\%$ for RMSE Rotation and Translation, respectively.
@@ 1 Introduction
Reassembly requires placing each element in its correct position and orientation to form the original shape as a whole – whether it be a 2D or 3D object. This ability is a form of spatial intelligence, which refers to the capacity to accurately perceive the visual-spatial environment and manipulate that perceived space [gardner2011frames]. This skill is typically evaluated through reassembly tasks, where individual components must be arranged and connected to form a coherent, functional whole—such as solving 2D jigsaw puzzles or assembling 3D structures with LEGO blocks.
Since the advent of the first puzzle solver [freeman1964apictorial], reassembly tasks have posed a significant challenge to the machine learning community due to their inherent combinatorial complexity. These challenges are further underscored by their wide range of applications, including genomics [marande2007mitochondrial], assistive technologies [yurteri2022obstacle], and molecular docking [corso2023diffdock].
@@
In recent years, AI techniques for reassembling objects have gained increasing traction in the heritage field, particularly in fresco reconstruction [villegas2021matchmakernet]. This growing application of AI is driven by the fundamental challenge archaeologists face in reconstructing the past. After extensive and painstaking work involving site surveys and excavations, they are often confronted with the daunting task of reassembling countless fragments of varying sizes, shapes, and appearances to recreate ancient artifacts or artworks. Depending on the complexity of the artifact, this reassembly process can span months, years, or even decades. In cases involving a vast number of pieces, the task may become nearly insurmountable, regardless of the skill of the experts. Despite recent advances, state-of-the-art methods for solving this task remain far from achieving a satisfactory accuracy [tsesmelis2025re]. The challenges persist not only in the 3D domain but also in the 2D space, where the reduced degrees of freedom do not significantly alleviate the complexity of the problem. Reduction to 2D has been used for a long time to aid in finding the solutions as many heritage objects, e.g. frescos, are planar on the dominant surface [funkhouser2011learning]. Traditional approaches use keypoints or patches [sagiroglu2006texture, nielsen2008solving, tsamoura2009automatic, derech2021solving] benefiting from optimization strategy allowing them to scale to complex irregular shapes, in contrast to deep learning approaches. Three primary factors contribute to the difficulty of fresco reassembly for deep learning approaches. First, the inherent complexity of irregular pieces and the diverse textures present in the fragments (Figure 1). Secondly, as a byproduct of the first, complex polygonal shapes are difficult to model in deep learning approaches, where comparisons need to be made across the points of one fragment edge to another fragment. Finally, deep learning methods typically require large-scale datasets to effectively learn meaningful patterns [sun2017revisiting], a requirement that is difficult to meet in most applications domains of reassembly.
@@
To address these challenges, we propose ReassembleNet, a scalable deep learning method that integrates a diffusion process with multi-modal feature fusion. Our approach enables scalability on irregular 2D shapes through keypoint selection, a crucial capability for highly fragmented structures that existing methods struggle to process [hosseini2023puzzlefusion]. We introduce a semi-synthetic dataset specifically designed to highlight the characteristics of fresco pieces. Through fine-tuning, we demonstrate that pre-training on a semi-synthetic dataset significantly enhances model performance on real-world frescos. We combine the keypoint selection with geometry and texture representation to constrain the matching problem modeled through an inter and intra-piece attention block. Finally, we wrap this model in a diffusion process to iteratively denoise the pose of pieces into the final locations.
@@ 2.1 Reassembly of Square-based Puzzles
Square-based jigsaws have been extensively studied in the literature, with early approaches using greedy methods. For instance, [pomeranz2011fully] introduced a three-phase greedy approach involving placement, segmentation, and shifting, relying on compatibility functions and estimation metrics. While effective for large puzzles, it requires predefined orientations and does not handle missing pieces. Similarly, [gallagher2012jigsaw] proposed a tree-based algorithm for puzzles with unknown orientations and locations, using Mahalanobis Gradient Compatibility (MGC) for boundary analysis. However, it struggles with color-based similarity and missing pieces. [paikin2015solving] presented a greedy strategy for puzzles without prior information on piece sizes or orientations, refining the initial configuration. It works with mixed or missing pieces but struggles with noisy images. Other methods, such as [jin2014jigsaw], combine edge and content similarity using MGC, but face challenges with large-scale puzzles. [sholomon2014generalized] used a genetic algorithm based on color similarity to solve puzzles with unknown locations and orientations.
In recent years, several deep learning approaches have been developed for puzzle reconstruction. One prominent direction leverages generative models to reconstruct a complete image from an unordered set of puzzle pieces [talon2022ganzzle, talon2025ganzzle++]. These methods aim to infer the global structure of the puzzle, effectively generating missing spatial relationships. However, they are not applicable when the input pieces are rotated, as they assume a fixed orientation. Another line of research explores diffusion-based processes to iteratively refine the placement of puzzle pieces step by step [giuliari2023positional, scarpellini2024diffassemble]. These approaches have demonstrated effectiveness in solving the oriented puzzle problem, where piece orientations are known. However, they are restricted to regular puzzle settings, where pieces conform to a structured grid. For real-world applications, square-based approaches are often impractical, as broken objects typically have irregular shapes, noisy visual content, and eroded or missing edges between pieces. A naive solution is to create a regular patch that encloses the broken object; however, this drastically hinders the effectiveness.
@@ 2.2 Reassembly of Irregular Pieces
Before the adoption of deep learning, various techniques were developed to address this problem. For instance, [sagiroglu2006texture] proposed an artifact reassembly method using inpainting, texture synthesis, and FFT-based image registration. Their approach aligns fragments by maximizing correlation through FFT shift theory, effectively handling damaged edges. Similarly, [nielsen2008solving] uses a divide-and-conquer strategy, grouping pieces based on texture and color similarity. The method utilizes RGB and HIS color spaces for color matching and employs co-occurrence matrices to extract texture features. In [tsamoura2009automatic], the authors introduced a computer-aided method consisting of four key steps. Their approach reconstructs fragmented images by identifying adjacent pieces, matching contours using a Smith-Waterman-based method with color similarity, refining alignment through Iterative Closest Point, and assembling the final image based on optimized alignment angles. More recently, [derech2021solving] proposed a patch-based optimization method for artifact reconstruction, leveraging color, gradients, and geometric transformations to match fragments.
In recent years, significant progress has been made in applying deep learning to handle irregular input shapes. The first work in this area is [sholomon2016dnn], which introduced DNN-Buddies, a deep neural network model for evaluating piece compatibility. In [le2019jigsawnet], the authors enhanced puzzle piece matching using CNNs and adaptive boosting. They proposed a multi-graph search algorithm to replace greedy strategies, enabling the handling of missing pieces and low-texture areas. [cao20242d] developed a classification network for evaluating the compatibility of fragment pairs and matching irregular puzzle pieces. However, their approach requires square-shaped pieces and cannot accommodate irregular outputs. PuzzleFusion [hosseinipuzzlefusion] is based on a diffusion model and treats puzzle pieces as simple polygons with a limited number of vertices. This method is designed to work only on toy problems, as it processes all keypoints of the input polygons. As a result, it becomes impractical for real-world datasets, where a large number of keypoints lead to exponential memory consumption. Finally, PairingNet [zhou2024pairingnet] is a learning-based image fragment pair-searching and -matching approach. Their method employs a graph-based network for feature extraction, a linear transformer-based module for fragment pair-searching, and a weighted fusion module for pair-matching. By formulating a similarity matrix to infer adjacent segments.
@@ 3 ReassembleNet Pipeline
In this paper, we focus on reassembling fragmented 2D objects where the input of our method is a set of images representing fragments with irregular shapes. Following [hosseini2023puzzlefusion], our method takes as input $m$ images of unordered pieces. We extract keypoints from the images by applying Harris Corner Detector [harris1988combined]. Points act as a helpful cue in the arrangement of the pieces as subsets of the points on each piece can be aligned to recover the overall piece pose.
@@
Next, we aim at extracting a multi-modal keypoint feature designed to capture both the geometric and textural properties of each fragment. This allows us to fully embed the shape and visual characteristics of the pieces, which are essential for proper alignment. Finally, we utilize a combination of attention blocks to model the relationships between the fragments. This solution helps the method better fit the pieces together by analyzing their spatial and contextual connections, ensuring that they are aligned in the correct configuration. The model is wrapped within a diffusion process to progressively learn the correct alignment in terms of translation ${\textbf{s}_{i}^{m}}\in~\mathbb{R}^{2}$ and rotation $\textbf{r}_{i}^{m}=[\cos(\theta_{i}^{m}),\sin(\theta_{i}^{m})]^{\top}$ [zhou2019continuity].
@@ 3.1 KeyPoint Feature Representation
Texture Features. We have also focused on extracting features specifically related to the textures of the pieces. In particular, we extracted global texture features that summarize the general characteristics of the element, as well as local texture features that capture more detailed, localized information. For global texture features, we leveraged a pre-trained ResNet18 [he2016deep] to capture high-level texture representations across the entire element. For local texture features, we adopted a similar approach using the same pre-trained ResNet18. Specifically, we extract 32×32 patches centered on each selected keypoint. These localized patches are processed through ResNet18, where we remove the final classification layer and utilize the penultimate layer for feature extraction. This combination of global and local texture features allows our model to capture both the overall texture characteristics of the element as well as more detailed texture variations, which are crucial for accurately solving reassembly tasks.
@@ 3.2 Keypoint Selector Module
Non-Learnable Algorithm. For the non-learnable strategy, we employ the heuristic Farthest Point Sampling (FPS) [GONZALEZ1985293] to perform an initial pre-selection of the $k$-points. FPS is a well-established technique in the field of 3D point cloud processing and geometric sampling. It iteratively selects the point that is furthest from the current set, ensuring maximum dispersion throughout the entire data space. This approach guarantees that the selected keypoints are evenly distributed, capturing the underlying structure and geometry of the input data effectively. By leveraging FPS, our method can robustly cover the entire image or spatial domain with a fixed number of representative points.
Learnable Algorithm. We frame the keypoint selection task as a graph sparsification problem [ijcai2024p891]. In our formulation, the set of keypoints ${K}^{m}$ corresponds to the set of vertices $V^{m}$ of a complete graph $G^{m}$ (i.e. fully connected). Graph sparsification is then defined as the process of selecting a subset of nodes, i.e., keypoints or edges from $G^{m}$ to produce a sparser graph $\hat{G}^{m}$. In other words, the elements of $\hat{G}^{m}$ are a subset of those in $G^{m}$.
@@
Graph Transformer. The projected features are then processed by a Graph Transformer [shi2020masked] to aggregate and refine the input information.
Pooling Layer. Finally, we apply a pooling layer [gao2019graph], where we select $k$ nodes from the original graph. The selection of nodes to drop is guided by a projection score computed against a learnable vector p. To ensure that gradients propagate into p, these projection scores also serve as gating values, allowing nodes with lower scores to retain fewer features. Defined as:
@@
where $D$ represents the feature matrix of the input graph $G^{m}$, $\hat{D}$ denotes the output feature matrix of the subset graph $\hat{G}^{m}$, $\|\cdot\|$ denotes the L2 norm, $\text{top-}k$ selects the top $k$ indices from a given input vector, $\odot$ represents element-wise multiplication, and $_{j}$ is an indexing operation that extracts slices at the indices specified by j. This operation involves only a point-wise projection and slicing of the original feature and adjacency matrices, thereby preserving sparsity.
@@ 3.3 Estimating Position and Rotation with Diffusion
We adopt Diffusion Probabilistic Models as defined in Denoising Diffusion Implicit Models (DDIM) [song2020denoising]. For each piece $m$, we apply an initial transformation, consisting of a translation $\textbf{s}^{m}$ and a rotation $\textbf{r}^{m}$, that is replicated identical for all its keypoints. To represent this initial transformation compactly, we define a concatenated vector for each keypoint $\textbf{x}_{i0}^{m}=[\textbf{s}_{i0}^{m^{\top}},\textbf{r}_{i0}^{m^{\top}}]^{\top}$, where $\textbf{s}^{m}_{i0}$ and $\textbf{r}^{m}_{i0}$ denote the initial translation and rotation applied to all keypoints of the piece. Since this transformation is applied globally, all keypoints within the same piece share these initial transformation parameters.
At training time, we iteratively add noise sampled from a Gaussian distribution $\mathcal{N}(0,I)$ to the poses of each keypoints (Forward Process). Following that, for the denoising process, we train ReassembleNet to reverse the noising process (Reverse Process) and predict the initial poses of all keypoints for every piece. We denote the initial poses of all the keypoints as $X_{0}=\{X_{0}^{m}\}_{m=1}^{M},$ where for each piece, $X^{m}_{0}=\{\textbf{x}^{m}_{i0}\}_{i=1}^{K},$ represents the set of initial keypoint poses. Additionally, we denote the predicted keypoint poses at timestep t1t-1 as $\hat{X}_{t-1}=\{\hat{X}_{t-1}^{m}\}_{m=1}^{M}$. For the final result at time $t=0$, we take the average polygon Euclidean center position and the polygon rotation. Further details on the diffusion process we use can be found in Supplementary Material C.
The Architecture. To process the task, we aim to capture both intra-piece information and inter-piece (i.e. global information between pieces). The key idea is to handle these two types of information separately but in parallel. To achieve this, we employ a combination of attention layers. Intra-piece information is processed using a sparse self-attention block [child2019generating], which helps the network focus on obtaining consistent positions and rotations, across different keypoints within the same piece. While, inter-piece information is handled using a standard self-attention block [vaswani2017attention] between pieces.
Losses. Following [ho2020denoising] and standard practice in Diffusion Models, we train ReassembleNet to directly predict $\hat{X}_{0}$ rather than $\hat{X}_{t-1}$. We employ two loss functions to reconstruct the initial pose of each piece’s keypoints.